第44回環論および表現論シンポジウム（2011年）第二報

第 44 回環論および表現論シンポジウムのプログラムが出来ましたのでお知らせいたします. また懇親会もあわせてご案内申し上げます. このシンポジウムは以下の研究費の援助の元に行われます.

日本学術振興会科学研究費補助金基盤研究 (B) 研究代表者：山形邦夫（東京農工大学）日本学術振興会科学研究費補助金基盤研究 (C) 研究代表者：馬場良始（大阪教育大学）研究代表者：池畑秀一（岡山大学）研究代表者：河田成人（大阪市立大学）研究代表者：小池寿俊（沖縄高専）研究代表者：西中恒和（岡山商科大学）

プログラム責任者 伊山 修（名古屋大学） e-mail: iyama@math.nagoya-u.ac.jp

記

日程： ２０１１年９月２５日（日）〜２７日（火）

場所： 岡山大学 理学部本館（１号館）２１・２２・２４番教室〒 700-8530 岡山市北区津島中３丁目１番１号（会場責任者：飛田明彦（埼玉大学））（補助責任者：池畑秀一（岡山大学））

懇親会： 日時：９月２６日（月）午後１９時００分〜２１時００分会場：岡山大学南福利施設ピーチユニオン４Ｆレストラン http://www.okayama-u.ac.jp/jp/tsushima j.html （地図の真ん中にあります）参加費： 5,000 円

懇親会参加希望者は９月１６日（金）までに池畑秀一 (ikehata@ems.okayama-u.ac.jp) までご連絡願います. 多数の方のご参加をお待ちしています.

９月２５日（日） 8:40–9:30 Dan Zacharia （Syracuse University）

On modules of finite complexity over selfinjective algebras I

9:40–10:30 Dong Yang On derived simple algebras

10:50–11:40 斉藤義久（東京大学）前射影多元環と量子群の結晶基底 I

13:00–13:40

21 教室阿部弘樹（筑波大学）

Reflection for selfinjective algebras 24 教室 Erik Darpö （名古屋大学）

The representation rings of the dihedral 2-groups

13:50–14:30

21 教室水野有哉（名古屋大学）

APR tilting modules and quiver mutations 24 教室小原大樹（東京理科大学）

Hochschild cohomology of quiver algebras defined by two cycles and a quantum-like relation

14:50–15:30

21 教室 Joseph Grant （名古屋大学）

Derived autoequivalences and braid relations 24 教室古谷貴彦・速水孝夫（東京理科大学・北海学園大学）

Hochschild cohomology of cluster-tilted algebras of types A_{n} and D_{n}

15:40–16:20

21 教室 Martin Herschend （名古屋大学）

n-representation infinite algebras 24 教室吉澤毅（岡山大学）

Subcategories of extension modules related to Serre subcategories

16:40–17:30 吉野雄二（岡山大学） Cohen-Macaulay 加群の表現およびその退化への入門 I

17:40–18:30 Jay A. Wood （Western Michigan University） Applications of finite Frobenius rings to algebraic coding theory I ９月２６日（月） 8:40–9:30 Dan Zacharia （Syracuse University）

On modules of finite complexity over selfinjective algebras II

9:40–10:30 Dong Yang Recollements generated by idempotents

10:50–11:40 斉藤義久（東京大学）前射影多元環と量子群の結晶基底 II

13:00–13:40

21 教室竹花靖彦（函館工業高専）

A generalization of costable torsion theory 24 教室源泰幸（京都大学）

A note on dimension of triangulated categories

13:50–14:30

21 教室木村真弓・浅芝秀人（静岡大学）

Quiver presentations of Grothendieck constructions 24 教室相原琢磨・高橋亮（千葉大学・信州大学）

Dimensions of derived categories

14:50–15:30

21 教室上山健太（静岡大学）

Graded Frobenius algebras and quantum Beilinson algebras 24 教室古賀寛尚・星野光男（筑波大学）

Weak Gorenstein dimension for modules and Gorenstein algebras

15:40–16:20

21 教室松岡学（四日市高校）

Polycyclic codes and sequential codes 24 教室平松直哉（呉高専)

On a degeneration problem for Cohen-Macaulay modules

16:40–17:30 吉野雄二（岡山大学） Cohen-Macaulay 加群の表現およびその退化への入門 II

17:40–18:30 Jay A. Wood （Western Michigan University） Applications of finite Frobenius rings to algebraic coding theory II

19:00– 懇親会９月２７日（火） 8:40–9:30 高橋篤史（大阪大学）

TBA

9:40–10:30 宮地兵衛（名古屋大学） Hidden Hecke Algebras and Duality

10:50–11:30

21 教室山浦浩太（名古屋大学） Tilting theory for stable graded module categories over graded self-injective algebras

24 教室中島規博（北海道大学） The Noetherian properties of the rings of differential operators on central 2-arrangements

12:50–13:30

21 教室長瀬潤・名倉誠（東京学芸大学・奈良高専）

Hom-orthogonal partial tilting modules for Dynkin quivers 24 教室岡崎亮太・柳川浩二（大阪大学・関西大学）

Alternative polarizations of Borel fixed ideals and Eliahou-Kervaire type resolution

13:40–14:20

21 教室 Laurent Demonet （名古屋大学） Categorification of cluster algebra structures of coordinate rings of simple Lie groups

24 教室松田一徳（名古屋大学） Weakly closed graph

14:40–15:20

21 教室木村嘉之（京都大学）

Quantum unipotent subgroup and dual canonical basis 24 教室大関一秀（明治大学）

Sharp bounds for Hilbert coefficients of parameters

15:30–16:20 本瀬香 The Example by Stephens